A total of $3700 was invested, part of it at 6% interest and the remainder at 11%. If the total yearly interest amounted to $330, how much was invested at each rate?
Answer by mananth(16946)

Part I 6.00% per annum ------------- Amount invested =x
Part II 11.00% per annum ------------ Amount invested = y
3700
Interest----- 330.00

Part I 6.00% per annum ---x
Part II 11.00% per annum ---y
Total investment
x + 1 y= 3700 -------------1
Interest on both investments
6.00% x + 11.00% y= 330
Multiply by 100
6 x + 11 y= 33000.00 --------2
Multiply (1) by -6
we get
-6 x -6 y= -22200.00
Add this to (2)
0 x 5 y= 10800
divide by 5
y = 2160 - x)
Part I 6.00% $ 1540
Part II 11.00% $ 2160

CHECK
1540 --------- 6.00% ------- 92.40
2160 ------------- 11.00% ------- 237.60
Total -------------------- 330.00

m.ananth@hotmail.ca